OpenStudy (idealist10):

How to solve for y: 4x^8-3x^5*y-3x^3*y-y^2+3x^6=0

OpenStudy (idealist10):

@hartnn

OpenStudy (idealist10):

\[4x^8-3x^5y-3x^3y-y^2+3x^6=0\]

OpenStudy (idealist10):

@zepdrix @ganeshie8 @amistre64 @Destinymasha @radar @waterineyes

OpenStudy (anonymous):

\[4x^8-3x^5y-3x^3y-y^2+3x^6=0 \]

OpenStudy (idealist10):

How to solve for y??

OpenStudy (anonymous):

I am also thinking for that.. :P

OpenStudy (anonymous):

I also need experts' advice here.. :) Nothing more I can help than ( @hartnn @ganeshie8 ) this..!! :P Can we get help here?

OpenStudy (anonymous):

zepdrix.. :)

zepdrix (zepdrix):

D:

OpenStudy (idealist10):

@hartnn

OpenStudy (anonymous):

Don't worry, they would be busy, wait for some time, they will surely check it out once.. :)

zepdrix (zepdrix):

Wolfram isn't showing any nice factored form. Are you sure this can be factored? :o

zepdrix (zepdrix):

Oh solve for y, my bad :3

zepdrix (zepdrix):

\[\Large\rm 4x^8+3x^6-3x^5y-3x^3y-y^2=0\]Multiply everything by -1, then start grouping some stuff,\[\Large\rm y^2+3x^3(x^2-1)y-x^6(4x^2+1)=0\]And then just throw it into the quadratic formula, yah? :O That should give you y in terms of x.

OpenStudy (idealist10):

So a=1, b=3x^3(x^2-1), c=-x^6(4x^2+1), and I plug these values into the quadratic formula, right?

hartnn (hartnn):

right. Just to make sure, whats the original question ?

OpenStudy (idealist10):

I got \[y^2+3x^3(x^2+1)y-x^6(4x^2+3)=0\]

OpenStudy (anonymous):

Solve for \(y\): \[4x^8-3x^5y-3x^3y-y^2+3x^6=0\]

OpenStudy (anonymous):

Yes, you are right.. @Idealist10

OpenStudy (anonymous):

zepdrix has done some typo or misinterpreting there.. Don't worry, you go ahead.. :) Good..!!

OpenStudy (idealist10):

I got it! Thanks!